prove : div cur l f =0


1
Expert's answer
2021-07-09T06:40:02-0400

We introduce the vector differential operator "\nabla" (“del”) as


"\nabla=\vec i \dfrac{\partial}{\partial x}+\vec j \dfrac{\partial}{\partial y}+\vec k \dfrac{\partial}{\partial z}"

Then for the vector field "\vec F"


"curl\vec F=\nabla \times\vec F"

"div\vec F=\nabla \cdot\vec F"

The cross product "\nabla \times\vec F"  is perpendicular to both "\nabla" and "\vec F."

Hence for any vector field "\vec F"

"\nabla\cdot(\nabla\times\vec F)=0"

Therefore for any vector field "\vec F"

"div(curl \vec F)=\nabla\cdot(\nabla\times\vec F)=0"